If an inventor of a new product charges $4 per unit, then her profit for producing and selling X units given by the function: P(x)= 2.25x-7000.
A. What is her profit if she sells 2000 units
B. What is her profit if she sells 5000 units
C. What is the break even point for sales

Found 2 solutions by richwmiller, palanisamy:
Answer by richwmiller(17219)

A)Plug in P(2000)=2.25(2000)-7000
b)Plug in P(5000)=2.25(5000)-7000
C)2.25x=7000 solve for x
x=7000/2.25

Answer by palanisamy(496)

The inventor of a new product charges $4 per unit.
Her profit for producing and selling x units given by the function:
P(x)= 2.25x-7000
A) When x = 2000, Profit P(2000) = (2.25)*2000-7000
= 4500-7000
= -2500
So there is a loss of 2500$
B)When x = 5000, Profit P(2000) = (2.25)*5000-7000
= 11250-7000
= 4250
So there is a profit of 4250$
C. What is the break even point for sales
To get the break even point we have P(x) = 0
2.25x-7000=0
2.25x = 7000
x = 7000/225
x = 31 nearly